# Separating Tongue from Thought: Activation Patching Reveals Language-Agnostic Concept Representations in Transformers

### v2

@@ Abstract
A central question in multilingual language modeling is whether large language models (LLMs) develop a universal concept representation, disentangled from specific languages. In this paper, we address this question by analyzing latent representations (latents) during a word translation task in transformer-based LLMs. We strategically extract latents from a source translation prompt and insert them into the forward pass on a target translation prompt. By doing so, we find that the output language is encoded in the latent at an earlier layer than the concept to be translated. Building on this insight, we conduct two key experiments. First, we demonstrate that we can change the concept without changing the language and vice versa through activation patching alone. Second, we show that patching with the mean over latents across different languages does not impair and instead improves the models’ performance in translating the concept. Our results provide evidence for the existence of language-agnostic concept representations within the investigated models. [^1] This work has been previously published under the title ”How Do Llamas Process Multilingual Text? A Latent Exploration through Activation Patching” at the ICML 2024 mechanistic interpretability workshop https://openreview.net/forum?id=0ku2hIm4BS.
Code for reproducing our experiments is available at https://github.com/Butanium/llm-lang-agnostic.
@@
Figure 1: For two given concepts, e.g., book and lemon, we construct multiple source prompts which translate BOOK, and a target prompt for translating from French to Chinese. Then we extract the residual stream of the last token of the word to be translated after some layer $j$ and all subsequent ones from the source prompts and insert the mean of each layer at the corresponding positions in the forward pass of the target prompt. The resulting next token probabilities will concentrate on the source concept in target language ($\text{{{book}}}^{\text{{{zh}}}}$, i.e., {CJK}UTF8gbsn柠檬) when patching at layers 0–15, on the target concept in target language ($\text{{{lemon}}}^{\text{{{zh}}}}$, {CJK}UTF8gbsn书) for layers 16–31
@@ 1 Introduction
The emergence of the field of mechanistic interpretability has led to the conception of powerful tools Carter et al. (2019); Nostalgebraist (2020); Schubert et al. (2020); Belrose et al. (2023); Cunningham et al. (2023); Kramár et al. (2024); Marks et al. (2024); O’Neill and Bui (2024); Tufanov et al. (2024) for the investigation of the inner workings of deep neural networks such as large language models (LLMs) Vaswani et al. (2017); Radford et al. (2019); Touvron et al. (2023) with the ultimate goal of reverse engineering the algorithms encoded in their weights. As a result, researchers today are often able to open up a “black box” neural network, and with near surgical precision pinpoint where a certain input-output behaviour comes from Wang et al. (2022); Conmy et al. (2023); Nanda et al. (2023); Zhong et al. (2024); Furuta et al. (2024).
This provides a unique opportunity to examine how multilingual concepts are represented and processed within LLMs, potentially revealing insights into language biases and concept formation. In particular, approaches based on activation patching Variengien and Winsor (2023); Ghandeharioun et al. (2024); Chen et al. (2024) in which activations are patched from one forward pass into another one while observing the output (c.f. Fig. 2) present a simple, yet effective way to inspect the representations learned and computations performed by LLMs.
Summary of contributions. In this work, we leverage activation patching to understand how LLMs process multilingual text, with a particular focus on whether they use a language-agnostic concept space, as theorized by Wendler et al. (2024). We focus on autoregressive, decoder-only transformers as they are the most common type of LLMs. In order to do so, we design multiple patching experiments leveraging pairs of translation prompts with differing expected predicted concept and language.
@@
    First, we perform an activation patching analysis of Llama 2 7B Touvron et al. (2023), patching at the last token position. We demonstrate that the model processes translation tasks by first resolving output language, then the concept to be translated.
@@
    Finally, in App. C we show that our observations generalize to a diverse set of transformer models varying in size, architecture, and training data, including Llama 2 70B, Llama 3 8B Dubey et al. (2024), Mistral 7B Jiang et al. (2023), Qwen 1.5 7B Bai et al. (2023) and Aya 23 8B Aryabumi et al. (2024).
Our results align with the theory for Llama 2 outlined by Wendler et al. (2024) and with previous results on BERT models Devlin et al. (2018) by Wu et al. (2019); Pires et al. (2019). However, our analysis goes beyond the observational approach using the logit lens by Wendler et al. (2024) and cross-lingual embedding similarity analysis by Wu et al. (2019) and Pires et al. (2019). Using activation patching, we perform causal interventions that allow us to draw stronger conclusions about the computations performed and representations used within the models.
@@ 2 Background
2.1 Transformers’s forward pass
When an autoregressive decoder-only transformer Vaswani et al. (2017); Touvron et al. (2023) with $L$ layers processes a sequence of input tokens $x_{1},\dots,x_{n}\in V$ from a vocabulary $V$, each token is initially transformed into a $d$-dimensional vector $h_{i}^{0}$ by an embedding layer. This first set of vector is the beginning of the residual stream. Then, for each token position $i$, the layer $j\in 1,\dots,L$ updates the residual stream the following way:
@@ 2.1 Transformers’s forward pass
2.2 Concepts
@@
2.3 Prompt design
@@
English: “lake” - Français: “lac" English: “cloud" - Français: “
@@
Importantly, whether the model correctly answers the prompt is determined by its next token prediction. For example above, the next token predicted should be “nu", the first token of “nuage". Thus, we can track $P(C^{\ell})$, i.e., the probability of the concept $C$ occurring in language $\ell$, by simply summing up the probabilities of all starting tokens of $w(C^{\ell})$ in the next-token distribution.
@@ 3.1 Experimental design
We start our analysis with an exploratory experiment on Llama 2 7B Touvron et al. (2023). We use 5-shots translation prompts to create paired source $S=\text{TP}(\ell^{\text{(in)}}_{S},\ell^{(\text{out})}_{S},C_{S})$ and target prompt $T=\text{TP}(\ell^{\text{(in)}}_{T},\ell^{(\text{out})}_{T},C_{T})$ datasets with different concept, input languages and output languages. If not mentioned otherwise, $\ell_{S}$ and $\ell_{T}$ refer to the output language of $S$ and $T$.
@@ 3.2 Results
Hypotheses. We are left with two hypotheses compatible with these results, depicted in Figure 4:
@@
Figure 4: Our hypothesis about how the forward pass could look like on our translation prompts. Every block consists of multiple transformer blocks and in between the blocks we denote the relevant content contained in the latents (in the residual stream). Because in hypothesis 2 concept and language cannot be disentangled one input-output language specific translation circuit per language pair is required.
@@ 4 Ruling out hypotheses
Further evidence experiment. In the experiments in Sec. 3 we did not observe source concept in target language. However, both H1 and H2 would allow for that to happen via patching in the right way. Therefore, in this experiment, instead of overwriting the residual stream at the last token of the prompt, we overwrite them at the last token of the word to be translated. Let $\rho_{S}$ and $\rho_{T}$ denote the position of that token in source and target prompt respectively. Since the concept information seems to enter via multiple layers (16-20) into the latent of the last token, we overwrite the latent corresponding to the token at position $\rho_{T}$ at layer $j$ and all subsequent ones. By patching in this way in both H1 and H2 we would expect to see large $P(C_{S}^{\ell_{T}})$.
Formally, we patch by setting $h_{\rho_{T}}^{(j)}(T)=h_{\rho_{S}}^{(j)}(S),\dots,h_{\rho_{T}}^{(L)}(T)=h_{\rho_{S}}^{(L)}(S)$.
@@
$h_{\rho_{T}}^{(\alpha)}(T)=\frac{1}{k}\sum_{i=1}^{k}h_{\rho_{S_{i}}}^{(\alpha)}(S_{i}),$
@@
A visualization of this argument can be seen in Figure 5
Figure 5: Illustration of two hypotheses on the latent representation of concepts across multiple languages:
H1: When given prompts like Deutsch: “Dorf" - Italiano: … Deutsch: “Buch" with different input languages, we expect to obtain a salient representation of the language-agnostic concept (e.g., a book, shown as the left vectors in orange) along with some task-irrelevant information (the right vectors in blue). Averaging these representations should yield a concept representation at least as salient as those from individual prompts.
H2: In this scenario, concept vectors differ due to language entanglement. As a result, we would not expect the mean vector to be as interpretable as those from single-language prompts.
Results. In the first experiment we use the same languages as above and in the second one we used de, nl, zh, es, ru as input and it, fi, es, ru, ko as output languages for the source prompts, and, fr to zh for the target prompt.
In Figure 6 we observe, that in both experiments we obtain very high probability for the source concept in the target language $P(C_{S}^{\textsc{{zh}}})$ from layers 0 to 15, i.e., exactly until the latents at the last token stop attending to the last concept-token.
@@
Additionally, Figure 6 (b) supports that we are in H1 and not in H2 because patching in the mean keeps $P(C_{S}^{\textsc{{zh}}})$ intact and even increases it. Therefore, instead of observing interference between the different language-entangled concepts as would have been predicted by H2, we observe a concept-denoising effect by averaging multiple language-agnostic concept representations which only makes sense under H1. Taking the mean over concept representations corresponding to different input languages seems to act like a majority voting mechanism resulting in an increase in $P(C_{S}^{\textsc{{zh}}})$.  [^3] Conversely, e.g., averaging over different translation prompt contexts but while keeping the input and output language fixed does not lead to an increase in $P(C_{S}^{\textsc{{zh}}})$ (see App. Figure 10 (b)).
@@
Figure 6: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use multiple source input languages de, nl, zh, es, ru and output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt in (b) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src" stands for source and “tgt" for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.

### v3

@@ Abstract
A central question in multilingual language modeling is whether large language models (LLMs) develop a universal concept representation, disentangled from specific languages. In this paper, we address this question by analyzing latent representations (latents) during a word-translation task in transformer-based LLMs. We strategically extract latents from a source translation prompt and insert them into the forward pass on a target translation prompt. By doing so, we find that the output language is encoded in the latent at an earlier layer than the concept to be translated. Building on this insight, we conduct two key experiments. First, we demonstrate that we can change the concept without changing the language and vice versa through activation patching alone. Second, we show that patching with the mean representation of a concept across different languages does not affect the models’ ability to translate it, but instead improves it. Finally, we generalize to multi-token generation and demonstrate that the model can generate natural language description of those mean representations. Our results provide evidence for the existence of language-agnostic concept representations within the investigated models. [^1] Code for reproducing our experiments is available at https://anonymous.4open.science/r/llm-lang-agnostic-1A75/
@@
Figure 1: For two given concepts, e.g., book and lemon, we construct multiple source prompts which translate BOOK, and a target prompt for translating from French to Chinese. Then we extract the residual stream of the last token of the word to be translated after some layer $j$ and all subsequent ones from the source prompts and insert the mean of each layer at the corresponding positions in the forward pass of the target prompt. The resulting next token probabilities will concentrate on the source concept in target language ($\text{{{book}}}^{\text{{{zh}}}}$, i.e., {CJK}UTF8gbsn书) when patching at layers 0–15, on the target concept in target language ($\text{{{lemon}}}^{\text{{{zh}}}}$, {CJK}UTF8gbsn柠檬) for layers 16–31.
@@ 1 Introduction
Most modern large language models (LLMs) are trained on massive corpora dominated by English text Touvron et al. (2023); Dubey et al. (2024); Radford et al. (2019); Brown et al. (2020); OpenAI (2023). Despite this imbalanced training, they achieve remarkable performance across multiple languages Shi et al. (2022), raising fundamental questions about how they process and represent multilingual information. Understanding these mechanisms is crucial not only to improve model performance, but also to identify potential biases and limitations in cross-lingual processing.
l question in multilingual language modeling is whether LLMs develop universal concept representations that transcend specific languages Wendler et al. (2024). For example, when a model processes the word "cat" in English and "chat" in French, does it map these to the same internal representation of the feline concept, or does it maintain separate language-specific representations? Several recent works hint at the existence of language-agnostic concept representations.
Wendler et al. (2024) found that for simple multilingual tasks independent of the input and output language intermediate decodings of concept representations using the logit lens Nostalgebraist (2020) decode to the English before they decode to the target language. Additionally, it has been long observed that instruction and safety tuning LLMs only on English data generalizes to other languages Li et al. (2024); Shaham et al. (2024). The presence of language-agnostic representation in the pretrained LLMs would provide an explanation for both of these behaviors.
This provides a unique opportunity for us to dig deeper and examine how multilingual concepts are represented and processed within LLMs, potentially revealing insights into language biases and concept formation. In particular, we are inspired by recent mechanistic interpretability approaches based on activation patching Variengien and Winsor (2023); Ghandeharioun et al. (2024); Chen et al. (2024a). These approaches are based on the idea of patching activations from one forward pass into another while observing the output (c.f. Fig. 2) and present a simple, yet effective way to inspect the representations learned and causally understand the computations performed by LLMs.
Summary of contributions. More specifically, in this work we aim to understand how transformers process and represent concepts during translation tasks, whether language and concept information can be manipulated independently in the model’s computations, and, whether models maintain separate language-specific concept representations or develop a shared conceptual space. To this end, we make the following contributions.
@@
    First, we perform an activation patching analysis of Llama 2 7B Touvron et al. (2023). We demonstrate that the model processes translation tasks by first resolving output language, then the concept to be translated.
@@
    We show that our observations generalize to a diverse set of transformer models varying in size, architecture, and training data, including Llama 2 70B, Llama 3 8B Dubey et al. (2024), Mistral 7B Jiang et al. (2023), Qwen 1.5 7B Bai et al. (2023), Aya 23 8B Aryabumi et al. (2024) and Gemma 2 2B Team et al. (2024).
  5. 5.
    Finally, to support our claim that mean representation are usable by the model in a autoregressive generation setting, we present a novel activation patching setup depicted in Figure 5 to show that a model can successfully write a definition of such a mean representation.
2 Related Work
LLMs have demonstrated remarkable capabilities in processing multilingual text across languages, with examples including encoder-only model like mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020a), and mT5 Xue et al. (2021) and decoder-only model like Aryabumi et al. (2024); Dubey et al. (2024). Studies on encoder-only models have shown that they develop language-agnostic representations, explaining their cross-lingual transfer capabilities. The methodology used was embedding similarity analysis Wu et al. (2019); Libovický et al. (2020); Muller et al. (2021); Conneau et al. (2020b) and probing methods Choenni and Shutova (2020); Pires et al. (2019).
Whiled decoder-only transformers are not primarily designed to develop contextual embedding, but rather for open-ended text generation, they also develop cross-lingual generalization, for example, when safety and instruction tuning Li et al. (2024); Chirkova and Nikoulina (2024). Mechanistic interpretability has led to powerful tools to analyze the language-agnosticity of these models. Using neuron analysis Stańczak et al. (2022); Chen et al. (2024b); Cao et al. (2024); Zeng et al. (2024); Tang et al. (2024) have shown that LLMs develop both language-agnostic and language-specific neurons. Wendler et al. (2024); Wu et al. (2024) use the logit lens Nostalgebraist (2020) to perform early decoding during the forward pass of both LLMs and VLMs and show that, no matter the language or input modality, the intermediate decodings concentrate on English before decoding to a specific language in the very last layers.
While the analyses via embedding similarity, probing and the logit lens give use valuable insight into the structure of the representation, they are not causal. For example, using logit lens insights from Wendler et al. (2024); Wu et al. (2024) does not allow to do counter factual intervention Anonymous (2024). Additionally, while the neuron level analysis studies the causal effects of the neurons, they do not study the representation themselves. Our work aims to fill this gap.
In order to do so, we draw inspiration from activation patching, which was introduced by Meng et al. (2022) and has been the main tool used to draw causal interpretation of LLM representations  Variengien and Winsor (2023); Geiger et al. (2022); Kramár et al. (2024). More recently, Ghandeharioun et al. (2024); Chen et al. (2024a) also show patching setups in which they can use the model itself to analyze its own model internal. Inspired by those methods, we developped two novel patching experiments supporting the language-agnosicity of LLMs representation.
@@ 3 Background
Transformers’s forward pass. When an autoregressive decoder-only transformer Vaswani et al. (2017); Touvron et al. (2023) with $L$ layers processes a sequence of input tokens $x_{1},\dots,x_{n}\in V$ from a vocabulary $V$, each token is initially transformed into a $d$-dimensional vector $h_{i}^{0}$ by an embedding layer. This first set of vector is the beginning of the residual stream. Then, for each token position $i$, the layer $j\in 1,\dots,L$ updates the residual stream the following way:
@@
Activation patching. Activation patching is a causal intervention technique that allows us to understand how different components of a neural network contribute to its output. The key idea is to run two forward passes through the network – one on a source input and one on a target input – and copy (or "patch") activations from specific positions and layers of the source forward pass into the target forward pass. By observing how these interventions affect the model’s output, we can understand what information is encoded in different parts of the network and how it is used.
More formally, given a source input $S$ and target input $T$, we can patch activations at position $i,i^{\prime}$ and layer $j$ by setting $h_{i}^{(j)}(T)=h_{i^{\prime}}^{(j)}(S)$ during the target forward pass, where $h_{i}^{(j)}$ represents the activation at position $i$ and layer $j$. The change in the model’s output distribution provides evidence about what information was contained in the patched activation.
@@
Prompt design. We use the same template as Wendler et al. (2024): For a given concept $C$, input language $\ell^{(\text{in})}$, and output language $\ell^{(\text{out})}$, we construct a few-shot translation prompt $\text{TP}(\ell^{(\text{in})},\ell^{(\text{out})},C)$. This prompt contains examples of single-word translations from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$, concluding with the model being tasked to translate $C$ from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$.
$\ell^{(\text{in})}$: “$C_{1}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{1}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C_{n}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{n}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “
We denote $\text{TP}^{\text{collect}}(\ell^{(\text{in})},\ell^{(\text{out})},C)$ the translation prompt ending at the last token of $C^{\ell^{(\text{in})}}$.
We expect that the last token of such prompts is where the model stores its latent representation of $C^{\ell^{(\text{in})}}$.
@@
English: “computer” - Français: “ordinateur” English: “ant” - Français: “fourmi” English: “cloud” - Français: “
@@
Importantly, whether the model correctly answers the prompt is determined by its next token prediction. For example, above, the next token predicted should be “nu", the first token of “nuage". Thus, we can track $P(C^{\ell})$, i.e., the probability of the concept $C$ occurring in language $\ell$, by simply summing up the probabilities of all starting tokens of $w(C^{\ell})$ in the next-token distribution.
@@ 4.1 Experimental design
We start our analysis with an exploratory experiment on Llama 2 7B Touvron et al. (2023). We use 5-shots translation prompts to create paired source $S=\text{TP}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$ and target prompt $T=\text{TP}(\ell^{(\text{in})}_{T},\ell^{(\text{out})}_{T},C_{T})$ datasets with different concept, input languages and output languages. If not mentioned otherwise, $\ell_{S}$ and $\ell_{T}$ refer to the output language of $S$ and $T$.
@@ 4.2 Results
Hypotheses. We are left with two hypotheses compatible with these results:
@@ 5 Ruling out hypotheses
Further evidence experiment. In the experiments in Sec. 4 we did not observe source concept in target language. However, both H1 and H2 would allow for that to happen via patching in the right way. Therefore, in this experiment, instead of overwriting the residual stream at the last token of the prompt, we overwrite them at the last token of the word to be translated. In order to do that, for the source prompt, we use $\text{TP}^{\text{collect}}$ instead of TP ($S=\text{TP}^{\text{collect}}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$). This means that we collect the activations at the last token of $C_{S}^{\ell_{S}^{(\text{in})}}$.
Let $\rho_{T}$ denote the position of that token in the target prompt. Since the concept information seems to enter via multiple layers (16-20) into the latent of the last token, we overwrite the latent corresponding to the token at position $\rho_{T}$ at layer $j$ and all subsequent ones. By patching in this way in both H1 and H2 we would expect to see large $P(C_{S}^{\ell_{T}})$.
Formally, we patch by setting $h_{\rho_{T}}^{(j)}(T)=h_{-1}^{(j)}(S),\dots,h_{\rho_{T}}^{(L)}(T)=h_{-1}^{(L)}(S)$.
@@
$h_{\rho_{T}}^{(\alpha)}(T)=\frac{1}{k}\sum_{i=1}^{k}h_{-1}^{(\alpha)}(S_{i}),$
@@
Results. In the first experiment, we use the same languages as above and in the second one we used de, nl, zh, es, ru as input and it, fi, es, ru, ko as output languages for the source prompts and fr to zh for the target prompt.
In Figure 4 we observe that in both experiments, we obtain very high probability for the source concept in the target language $P(C_{S}^{\textsc{{zh}}})$ from layers 0 to 15, i.e., exactly until the latents at the last token stop attending to the last concept-token.
@@
Additionally, Figure 4 (b) supports that we are in H1 and not in H2 because patching in the mean keeps $P(C_{S}^{\textsc{{zh}}})$ intact and even increases it. Therefore, instead of observing interference between the different language-entangled concepts as would have been predicted by H2, we observe a concept-denoising effect by averaging multiple language-agnostic concept representations which only makes sense under H1. Taking the mean over concept representations corresponding to different input languages seems to act like a majority voting mechanism resulting in an increase in $P(C_{S}^{\textsc{{zh}}})$.  [^3] Conversely, e.g., averaging over different translation prompt contexts but while keeping the input and output language fixed does not lead to an increase in $P(C_{S}^{\textsc{{zh}}})$ (see App. Figure 11,12 (b)).
@@
Figure 4: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use multiple source input languages de, nl, zh, es, ru and source output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt in (b) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src" stands for source and “tgt" for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
@@
6 Generating descriptions for latents
We just showed that LLMs can effectively translate concept representations averaged across multiple languages, providing evidence for language-agnostic concept representations. In this section, we explore whether these mean representations, which theoretically capture language-independent concepts, can be described by the model in natural language as effectively as concepts expressed in a single language.
6.1 Definition prompt
In order to do that, we introduce a new prompt template that tasks the model to describe a concept in natural language. More precisely, given a concept $C$ and a language $\ell$, we construct a few shot definition prompt $\text{DP}(\ell,C)$ of the form:
$C^{\ell}_{1}$” : “$D^{\ell}_{1}$ $C^{\ell}$” : “
where $C^{\ell}_{1},\dots,C^{\ell}_{n}$ are concepts in language $\ell$ and $D^{\ell}_{1},\dots,D^{\ell}_{n}$ are their descriptions in language $\ell$. We denote $\text{DP}^{\text{collect}}(\ell,C)$ the prompt template that ends at the last token of $C^{\ell}$.
For each language, we constructed a definition dataset using the first concept in BabelNet Navigli et al. (2021) associated with each of the 200 picturable words from the Basic English word list from Wikipedia [^4] https://en.wiktionary.org/wiki/Appendix:Basic_English_word_list#Things_-_200_picturable_words. For each concept, BabelNet provides several definitions in different languages.
6.2 Patching setup
For languages $\ell_{S}^{1},\dots,\ell_{S}^{n}$ and $\ell_{T}$, and concepts $C_{S}\neq C_{T}$, we construct a target prompt $T=\text{DP}(\ell_{T},C_{T})$ and two sets $\mathcal{S}$ of source prompts:
  • From translations: for each language $\ell_{S}^{i}$ we pick an input language $\ell_{(\text{in})}^{i}$ and choose
    $\mathcal{S}_{\text{trans}}=\{\text{TP}^{\text{collect}}(\ell_{(\text{in})}^{i},\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
  • From definitions: we choose
    $\mathcal{S}_{\text{def}}=\{\text{DP}^{\text{collect}}(\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
Then, to generate a definition of $C_{S}$ in language $\ell_{T}$, for all layers, we patch the latents of the last token averaged over the source prompts from $\mathcal{S}$ to the last token of $C_{T}$ in the target prompt and let the model generate the definition as depicted in Figure 5. More formally, we patch by setting
$h_{\rho_{T}}^{(j)}(T)=\frac{1}{n}\sum_{i=1}^{n}h_{-1}^{(j)}(S_{i})$
for $j\in\{1,\dots,m\}$ and $\rho_{T}$ the position of the last token of $C_{T}$ in the target prompt.
Figure 5: Illustration of the patching setup for the definition prompt experiment. We patch the latents of the last token of the source prompts from $\mathcal{S}$ to the last token of $C_{T}$ in the target prompt.
6.3 Experiment
To compare the quality of the definitions of $C_{S}$ generated by the model in our experiment, we use the following baselines:
  • Ground truth: We use a random definition from BabelNet.
  • Prompting: We use the definition generated by the model when prompted with $\text{DP}(\ell_{T},C_{S})$.
  • Word patching: We replace $C_{T}^{\ell_{T}}$ with $C_{S}^{\ell_{S}^{j}}$ for a random $j\in\{1,\dots,n\}$ and let the model generate the definition. In this setting, the model is tasked to generate a definition of a word in a language different from the target prompt’s language.
  • Repeat word: We use $C_{S}^{\ell_{T}}$ as the definition.
To evaluate the different definitions, we use the sentence-transformers library Reimers and Gurevych (2019) to run paraphrase-multilingual-mpnet-base-v2555https://huggingface.co/sentence-transformers/paraphrase-multilingual-mpnet-base-v2, a semantic similarity model distilled from Song et al. (2020) using the methods from Reimers and Gurevych (2020). To assess the quality of the generated definition, we compute the similarity score between embeddings of the definition and the mean embedding of the ground truth definitions [^6] To be able to compare the score of the generated definition with the ground truth baseline, we compute its similarity score with the mean embedding of the other ground truth definitions - excluding the one that was randomly chosen as the ground truth baseline.
6.4 Results
Figure 6: Mean similarity between the definition and the mean embedding of the ground truth definitions, as well as the mean similarity between the definition embedding and the embeddings of the definitions of the other concepts in the dataset. The results are presented for three target languages: English (with source languages Italian, Finnish, Spanish, Korean, and input languages for source translations: German, Dutch, Chinese, Russian), Chinese (using the same languages), and French (with source languages Korean, Japanese, Estonian, Finnish and English as input language for source translations). We report means and 95% Gaussian confidence intervals computed over a dataset of various sizes [^8] For some concepts and languages, BabelNet does not provide any definition..
We report the mean similarity score between the definition and the mean embedding of the ground truth definitions for LLama 2 7B in Figure 8. To give an idea of the scale of the similarity scores, we also report the mean similarity between the definition embedding and the embeddings of the definitions of the other concepts in the dataset.
We find that patching concept representations from one language to another allows the model to generate high-quality definitions, comparable to or better than direct prompting. The fact that patching mean representations across multiple source languages leads to slightly better results suggests that the model’s concept representations are indeed language-agnostic – if they were language-specific, averaging across languages would likely degrade performance. This aligns with our translation experiment findings and adds another perspective on how LLMs process multilingual information.
Additionally, the comparable performance between patching from definitions and translations indicates that the model builds similar concept representations regardless of whether it processes a translation or definition prompt. This suggests a unified internal representation of concepts that generalizes across different types of language tasks.
Other models. In Appendix C.3 we show that those findings generalize to other models and languages.